Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2015-02-16

Observers : Laurent Le Guillou, Sylvain Baumont, Ayan Mitra

Support team : Nicolas Chotard, Pierre Francois Leget 

TAT : L. Galbany, M. Fraser


We did only follow-up observations, with EFOSC.


Clear sky, low wind, below 50% humidity.

Photometric night.


(times are in UT)


CALIBRATIONS :

 - Evening : EFOSC_bias, EFOSC_FlatSpec, EFOSC_SpecArc

          Again, the EFOSC rotator warmed up => A few BV SkyFlats.

 - Morning : EFOSC_SkyFlat: zi


SCIENCE OBSERVATIONS :


 - 00:54 L745a  spectro-photometric standard   g11+g13+g16, airmass 1.1


 - 01:06 seeing ~ 1.6 , humidity 60% , temperature 15º 


 - 01:28 PESSTOESO154 g16 2x2700s


 - 03:21 OGLE-2015-SN-043 g13 3x2700s 

       (good seeing, but there were optical aberrations, PSF seems not so good)


 - 05:42  image analysis


 - 05:54  L745a  spectro-photometric standard   g11+g13+g16, airmass 1.5 

        humidity 40%, seeing 0.6", temperature 17º


 - 06:24  LSQ15kp  g11 2x1800s, g16 2x1800s 


 - 08:48  PS1-15at  g13 1200s

        => SN-II, +50days, z ~ 0.03


 - 09:15 EG274  g11+g13+g16



CONCLUSIONS :


 PS1-15at  : SN-II, old (30 to 60 days), z ~ 0.03


     

Comments